OpenStudy (anonymous):

Using the first three terms of the Maclaurin's series, what is the value of cos(Ï€/2)?

OpenStudy (anonymous):

\[\cos (\pi/2) =1 - (\pi/2)^{2}/2! + (\pi/2)^{4}/4!\]